(50.0x2.5)-(11x5)
using significant figures
1
Expert's answer
2020-01-21T06:57:31-0500

50.0*2.5 = 125.0

we will round up 125.0 to 130 as 2.5 has only two significant figures

11*5 = 55

but we will round off this to 60 as 5 has only one significant figure

130-60=70

hence 70 will be the final answer as it contains only one significant figure